The sides of a triangle are x cm, (x - 4) and (x + 2) cm long respectively. If the permeter is 46cm, find the lenght of each side.
Answer by jim_thompson5910(35256)

x+x-4+x+2=46

3x-2=46

3x=46+2

3x=48

x=(48)/(3)

x=16

x - 4 = 16-4 = 12

x + 2 = 16 + 2 = 18

So the three sides are 16 cm, 12 cm and 18 cm.